# The Human Labour of Data Work: Capturing Cultural Diversity through World Wide Dishes

### v2

@@ 3. Placing World Wide Dishes in Historical Context
As described in Section 2, there are many concerns with biases in large-scale machine learning models which can be traced to problems in the underlying training and evaluation datasets. In this section, we explore the issue with long-tailed data distributions, where socially marginalised perspectives are disproportionately under-represented, and their relationship to bias in ML. We discuss how historic and systemic factors have resulted in these long-tailed distributions. As a result, ML models struggle to learn meaningful patterns or produce accurate, fair results for socially marginalised groups because long-tailed data comprise a negligible portion of the training instances (Zhang et al., 2023; Massiceti et al., 2021).
@@
The path for data to be included on the “Internet” is not equitable. Data from socially marginalised groups are less likely to be preserved in a digitised format and available on the Internet (Noble, 2018). Additionally, these data are systematically screened out of training datasets by state-of-the-art filtering models, such as CLIP (Radford et al., 2021), which are biased towards including data from and about Western cultures (Hong et al., 2024). Researchers use filtering methods such as CLIP to curate datasets in the first place to improve model quality by removing noisy, irrelevant, or potentially harmful content, aiming to optimise the relevance and accuracy of the data for the model’s intended use (Fang et al., 2023). However, these filtering methods often encode and amplify existing biases, leading to datasets that inadvertently exclude non-Western cultures and marginalised groups. Further, these filters are most likely to fail when these underrepresented groups are present in NSFW, racist, or stereotypical content (Birhane et al., 2021). This inconsistent filtering thus enhances and reinforces disparities in representation and increases concerns about the long-tailed nature of the datasets.
@@
The root causes of the long-tailed nature of data about socially marginalised groups are intimately intertwined with social, political, and economic histories. In other words, the systemic inclusion from—and misrepresentative nature of—datasets is not solely a product of technical failure. Rather, it is a manifestation of societal values contributing to choices made about whose cultures are recorded and whose are erased (Bowker, 2000; Cheney-Lippold, 2017; Brubaker and Hayes, 2011; Benthall and Haynes, 2019; Benjamin, 2019). For example, the exclusion and flattening of unique cultures from the African continent in digital media can be traced back to colonial regimes that sought to erase and thereby dehumanise the people living on the continent (Faloyin, 2022). During colonial times, power structures meant that the historical record was kept by Europeans colonising the continent (Bowker, 2000). In an attempt to justify their actions, they systematically flattened and dehumanised the lived experiences of those they sought to subjugate on the African continent. Colonial classification regimes flatten important cultural and geographic differences (Das and Semaan, 2022; Prabhakaran et al., 2022; Das et al., 2021). Borders were drawn without regard for existing complex social and ethnic groups, and the subsequent oppression and power dynamics involved in recording history lent themselves to a systematic erasure of distinct experiences (Bowker, 2000). More recent depictions of Africa in the international media and popular culture have made little attempt to capture its deeply complex and rich landscape (Faloyin, 2022).
@@
Acknowledging the systemic and historical reasons for this type of flattening, WWD explicitly seeks to counteract it through on-the-ground community consultation and collaboration. The granularity required to capture a region’s knowledge is most likely best understood through extensive consultation with communities with lived experiences in those regions. For example, borders may not provide the best guide for demarcating cultural boundaries, and a representation of “Kenya” or “Nigeria” may not be granular enough when we consider the distinct cultural groups within these borders. Some borders may also be misleading: for example, the Semliki River has changed course numerous times in the past few decades, meaning that some cultural groups have found themselves flipping between national identities of Uganda and the Democratic Republic of Congo, depending on a naturally evolving geographical marker (Faloyin, 2022).
@@
We present the African continent as an example to explain the deep-rooted causes of and potential for the type of cultural erasure that exists across the globe. WWD and the infrastructure developed therein seeks to counteract cultural erasure stemming from limitations in web-scraped datasets by encouraging data submissions from around the world in a granular manner, going beyond national representation but explicitly requesting fine-grained regional details about dishes and their associated cultural customs. We focused on food as a domain for this initiative because food is deeply intertwined with a region’s history, geography, and cultural practices. Preserving accurate and granular information about food is a crucial element in recording information about culture. For example, a large-scale dataset of dishes maps one singular dish called “pottage” onto eight different African countries, when in fact each of these countries is home to a specific, discrete version of the dish with unique ingredients (Winata et al., 2024). This kind of generalisation demonstrates how cultural nuance is lost when datasets fail to capture regional distinctions, reinforcing the need for granular approaches such as WWD (Magomere et al., 2025). In the coming sections, we expand on the dataset we created to counteract the flattening of cultural representation and the process of creating this dataset alongside community members.
@@ 4.1. The World Wide Dishes dataset
Table 1. Data collected through the World Wide Dishes contribution form
@@
The WWD dataset is the product of a bottom-up, community-led crowdsourcing effort to collect text, and when possible, image, data about regional dishes. We define bottom-up, community-led crowdsourcing efforts as those allowing any community member to engage and making space for community contributors to shape parts of the research process. The Core Organisers behind WWD launched a concerted effort to encourage data contributions throughout April 2024. At the end of April 2024, the WWD dataset contained 765 dishes with detailed metadata. This included 372 images verified by Community Contributors (301 images with a Creative Commons or royalty-free licence, and 71 user-uploaded images (shared with permission)). These dishes are associated with 106 countries (see Table E.2 for breakdown of regional coverage), with dish names in 131 languages. Community Contributors contributed information about a variety of dishes ranging from main dishes to desserts (see breakdown of dish types in Table E.3). Notably, 101 of these dishes share more than one region of association. Dishes in the WWD are accompanied by detailed metadata (see Table 1 for metadata information). Many of these metadata fields (e.g., time of day eaten, utensils used) were cultural aspects of dishes that the Core Organisers would not have known to capture, if not for direct input from community members. In the following subsection, we give an overview of the stakeholders involved in the development of WWD and a description of how they interacted with each other and with the overall project.
@@ 4.2. Stakeholders in the World Wide Dishes System
Figure 1. Stakeholders in World Wide Dishes. There are three stakeholder groups in WWD: Community Contributors, Community Ambassadors, and Core Organisers. This figure illustrates the overlapping stakeholder roles in the project. All Core Organisers were also Community Ambassadors and Community Contributors, while all Community Ambassadors were also Community Contributors. Responsibilities were additive, meaning Core Organisers took on all tasks assigned to Community Ambassadors and Community Contributors, and Community Ambassadors took on all tasks assigned to Community Contributors. Feedback flowed freely from Community Contributors to both Community Ambassadors and Core Organisers, who then implemented changes to the data collection platform based on this input.
WWD is a bottom-up, community-led dataset creation initiative that relied on three key stakeholder groups: Community Contributors (CCs), Community Ambassadors (CAs), and Core Organisers (COs). (Figure 1) The stakeholder groups are not mutually exclusive; for example, all COs also served as CAs. CAs were members of the regions that they oversaw (e.g., they were from and maintained close ties with or currently lived in the geographic region). Below we describe the responsibilities of each stakeholder group and how they interacted with each other.
@@
Core Organisers (COs): In addition to the responsibilities of CCs and CAs, COs coordinated and handled the administration of the WWD project. They began the initiative to build WWD, and developed and iterated the data collection process, front-end WWD project website, and back-end database, as well as the data processing pipeline in accordance with community feedback. COs ensured accessibility was a top priority and integrated feedback from the community continuously to achieve this goal. Additionally, they helped to review community contributions for dishes from their regions. All COs also served as CCs and CAs for their home regions.
Together, the three stakeholder groups made collecting fine-grained regional dish data possible. The CAs played a crucial role in mediating interactions between CCs and the overall research project. Because CAs were members of the communities for which they served as ambassadors, they had intimate knowledge of the potential barriers to participation that CCs might face. For example, when the data collection instrument for WWD was being developed, CAs raised the need to build a web interface that was compatible with the 2G cellular networks that were the norm for Internet connectivity in their home regions. In the following section, we describe the technical system architecture through which data contributions were made and then conclude this section with an overview of how community members interacted with and shaped the design of this system.
@@ 4.3.1. Data collection
Data collection occurred through a web-based form accessible on our website (see Figure B.1). On the web-based form, CCs were first provided with information about the project and a consent form. If they provided consent and indicated that they were over 18, they were then taken to the data collection form and asked to provide information about their chosen dish. As part of this form, CCs were asked to upload a non-generated image of food from personal records, indicate the dish name in the local language with a translation or phonetic equivalent to English when possible, and provide information about the customs associated with the dish. Please see Table 1 above for a full list of the fields CCs were asked to provide.
For the WWD data collection process to be useful, contributors had to be recruited and actively engaged. To that end, the COs and CAs worked together to ensure the data collection system was easily shareable via digital social networks to support multiple browsers, and also had interactive elements to amplify CC excitement and motivation. Recruitment involved the use of many social networks, including through existing communities such as the AYA Project, [^6] https://aya.for.ai/ Masakhane, [^7] https://www.masakhane.io/ the Deep Learning IndabaX network, [^8] https://deeplearningindaba.com/2024/indabax/ AI Saturdays Lagos, [^9] https://aisaturdayslagos.github.io/ and OLS [^10] https://we-are-ols.org/. Posts were also placed on a mailing list to encourage engagement.
Engagement was explicitly encouraged through the promotion of a submission leaderboard (see Figure 3), which displayed the number of dish submissions and contributors received from each region. This was done in an attempt to gamify the experience and build excitement and fun. The leaderboard was promoted during recruitment outreach.
Figure 3. Screenshot of the World Wide Dishes Learderboard showing the top 10 countries by number of dishes contributed, the total number of dishes contributed, and the total number of contributors per country.
@@ 4.3.2. Data storage
The collected data is stored in two sections: the dish data, which has been made public and is completely anonymised, and the personal data, which has been collected and stored according to the terms agreed by our ethics review board. We collected names, ages, and national identities, and also accepted anonymous submissions. Names were only used when explicit consent was given in order to acknowledge contributions publicly. Age and national identity data were required to help us understand the demographics of people submitting to WWD. Approval for data collection and the subsequent research study was obtained from the Departmental Research Ethics Committees of the Computer Science Department at the University of Oxford (reference: CS_C1A_24_004). Notably, the ethics review required the collection of age data, as CCs had to be over the age of 18 to participate; age data was therefore collected even if the CC otherwise submitted dish data anonymously.
@@ 4.3.3. Data processing
    Inconsistencies in submitted data were handled by consulting CAs where possible, and CAs also consulted with CCs as needed. However, we aimed to collect lived experiences around the world and so we did not make any efforts to police a ‘ground truth’ for each dish, or assign an ‘origin’ or ‘authenticity’ to any dish. Multiple nations or cultural groups can share the same dishes, for a wide variety of reasons including historical trade routes, war and occupation, the redrafting of political borders without regard to cultural borders, and migration patterns. Our task was not to make judgements about cultural ‘ownership’ of the submitted dishes. Instead, the data collection asked participants to offer dishes from their own lived experiences and personal backgrounds, so we expected to see similarities across borders, as well as some reasonable regional and preference variations of the ingredients of dishes within and across borders.
@@ 4.3.4. How stakeholders interacted with and shaped the design of the WWD technical system architecture
All three stakeholder groups (CCs, CAs, and COs) interacted with and shaped the technical system through which data contributions were made. Stakeholders surfaced additional metadata fields that were important to understanding the traditions of their local dishes during focus groups prior to data collection, and the COs then adjusted the data collection form to include these fields. By guiding the questions asked on the data collection form, community members shaped both the technical data collection system and the resulting WWD dataset. Moreover, COs and CAs who had intimate knowledge of potential barriers to participation for their home regions directly shaped the technical data collection system by advocating for a web interface that was mobile-first and compatible with low-bandwidth cellular networks. CAs provided valuable insight during data collection as to sub-regions and cultures that were under-represented during data collection and COs and CAs responded by introducing extra efforts to reach these regions.
@@ 5. Methods and Analysis
Our study aims to map out the process of building a dataset of cultural objects—specifically food—through a bottom-up, community-led approach. In this section, we provide a clear timeline for developing WWD, detail our data collection and reflective practices, and explain our analysis methodology.
@@ 5.1. Timeline and project context
Planning, building, and deploying the system that resulted in the WWD dataset was a process that occurred over 12 months. In October 2023, the first author [^11] While the first and second author are joint first authors on this manuscript, for the remainder of the paper we will refer to them as the first and second author, respectively, indicating the order in which their names appear in the author list. We do this for simplicity as the first and second author had different positionalities. initiated the WWD project. From November to December 2023, the team expanded to include three more COs, with four more joining in March 2024 (see Table 2 for details about the full team list). The final addition of COs occurred in April 2025 in order to include two CAs in honour of their exceptional commitment to the project. Between January and February 2024, the CO team developed the data collection instrument and website. The WWDwebsite went live in February 2024. Inspired by the AYA initiative’s use of community ambassadors (Singh et al., 2024), the COs recruited and onboarded CAs throughout March 2024 to support community engagement. In April 2024, the COs and CAs conducted a month-long campaign for data contributions, resulting in 765 dishes submitted from 106 regions. The research activities conducted as part of the development of WWD were approved by the first author’s IRB.

### v3

@@ 3. Placing World Wide Dishes in Historical Context
As described in Section 2, there are many concerns with biases in large-scale machine learning models which can be traced to problems in the underlying training and evaluation datasets. In this section, we explore the issue with long-tailed data distributions, where socially marginalised perspectives are disproportionately underrepresented, and their relationship to bias in ML. We discuss how historic and systemic factors have resulted in these long-tailed distributions. As a result, ML models struggle to learn meaningful patterns or produce accurate, fair results for socially marginalised groups because long-tailed data comprise a negligible portion of the training instances (Zhang et al., 2023; Massiceti et al., 2021).
@@
The path for data to be included on the “Internet” is not equitable. Data from socially marginalised groups are less likely to be preserved in a digitised format and available on the Internet (Noble, 2018). Additionally, these data are systematically screened out of training datasets by state-of-the-art filtering models, such as CLIP (Radford et al., 2021), which are biased towards including data from and about Western cultures (Hong et al., 2024). Researchers use filtering methods such as CLIP to curate datasets to improve model quality by removing noisy, irrelevant, or potentially harmful content, aiming to optimise the relevance and accuracy of the data for the model’s intended use (Fang et al., 2023). However, these filtering methods often encode and amplify existing biases, leading to datasets that inadvertently exclude non-Western cultures and marginalised groups (Qadri et al., 2025). Ironically, while these filters are designed to protect against inappropriate or harmful content, they sometimes fail to remove problematic material such as foul language, racism, or harmful stereotypes (Birhane et al., 2021). This inconsistent filtering reinforces disparities in representation and increases concerns about the long-tailed nature of the datasets.
@@
The root causes of the long-tailed nature of data about socially marginalised groups are intimately intertwined with social, political, and economic histories. In other words, the systemic inclusion from—and misrepresentative nature of—datasets is not solely a product of technical failure. Rather, it is a manifestation of societal values contributing to choices made about whose cultures are recorded and whose are erased (Bowker, 2000; Cheney-Lippold, 2017; Brubaker and Hayes, 2011; Benthall and Haynes, 2019; Benjamin, 2019). For example, the exclusion and flattening of unique cultures from the African continent in digital media can be traced back to colonial regimes that sought to erase and thereby dehumanise the people living on the continent (Faloyin, 2022). During colonial times, power structures meant that the historical record was kept by Europeans colonising the continent (Bowker, 2000). In an attempt to justify their actions, they systematically flattened and dehumanised the lived experiences of those they sought to subjugate. Colonial classification regimes flatten important cultural and geographic differences (Das and Semaan, 2022; Prabhakaran et al., 2022; Das et al., 2021). Borders were drawn without regard for existing complex social and ethnic groups, and the subsequent oppression and power dynamics involved in recording history lent themselves to a systematic erasure of distinct experiences (Bowker, 2000). More recent depictions of Africa in the international media and popular culture have made little attempt to capture its deeply complex and rich landscape (Faloyin, 2022).
@@
Acknowledging the systemic and historical reasons for this type of flattening, WWD explicitly seeks to counteract it through on-the-ground community consultation and collaboration. The granularity required to capture a region’s knowledge is most likely best understood through extensive consultation with communities with lived experiences in those regions. For example, borders may not provide the best guide for demarcating cultural boundaries, and a representation of “Kenya” or “Nigeria” may not be granular enough when we consider the distinct cultural groups within these borders. Some borders may also be misleading: for example, the Semliki River has changed course numerous times over the past few decades, meaning that some cultural groups have found themselves flipping between national identities of Uganda and the Democratic Republic of Congo, depending on a naturally evolving geographical marker (Faloyin, 2022).
@@
We present the African continent as an example to explain the deep-rooted causes of and potential for the type of cultural erasure that exists across the globe. WWD and the infrastructure developed therein seeks to counteract cultural erasure stemming from limitations in web-scraped datasets by encouraging data submissions from around the world in a granular manner, going beyond national representation by explicitly requesting fine-grained regional details about dishes and their associated cultural customs. We focused on food because it is deeply intertwined with a region’s history, geography, and cultural practices. Preserving accurate and granular information about food is a crucial element in recording information about culture. For example, a large-scale dataset of dishes maps one singular dish called “pottage” onto eight different African countries, when in fact each of these countries is home to a specific, discrete version of the dish with unique ingredients (Winata et al., 2024). This kind of generalisation demonstrates how cultural nuance is lost when datasets fail to capture regional distinctions, reinforcing the need for granular approaches such as WWD (Magomere et al., 2025). In the coming sections, we expand on the dataset we created to counteract the flattening of cultural representation and the process of creating this dataset alongside community members.
@@ 4.1. The World Wide Dishes dataset
Table 1. Data collected through the World Wide Dishes contribution form.
@@
The WWD dataset is the product of a bottom-up, community-led crowdsourcing effort to collect text and, when possible, image data about regional dishes. The Core Organisers (see Section 4.2) behind WWD launched a concerted effort to encourage data contributions throughout April 2024. At the end of April 2024, the WWD dataset contained 765 dishes with detailed metadata. This included 372 images verified by Community Contributors (301 images with a Creative Commons or royalty-free licence, and 71 user-uploaded images (shared with permission)). These dishes are associated with 106 countries (see Table E.2 for breakdown of regional coverage), with dish names in 131 languages. Community Contributors contributed information about a variety of dishes ranging from main dishes to desserts (see breakdown of dish types in Table E.3). Notably, 101 of these dishes share more than one region of association. Dishes in WWD are accompanied by detailed metadata (see Table 1 for metadata information). Many of these metadata fields (e.g., time of day eaten, utensils used) are cultural aspects of dishes that the Core Organisers would not have known to capture, if not for direct input from community members. In the following subsection, we give an overview of the stakeholders involved in the development of WWD and a description of how they interacted with each other and with the overall project.
@@ 4.2. Stakeholders in the World Wide Dishes System
Figure 1. Stakeholders in World Wide Dishes. There are three stakeholder groups in WWD: Community Contributors, Community Ambassadors, and Core Organisers. This figure illustrates the overlapping stakeholder roles in the project.
WWD is a bottom-up, community-led dataset creation initiative that relied on three key stakeholder groups: Community Contributors (CCs), Community Ambassadors (CAs), and Core Organisers (COs) (Figure 1). The stakeholder groups are not mutually exclusive; for example, all COs also served as CAs. CAs were members of the regions that they oversaw (e.g., they were from and maintained close ties with or currently lived in the geographic region). Below we describe the responsibilities of each stakeholder group and how they interacted with each other.
@@
Core Organisers (COs): In addition to the responsibilities of CCs and CAs, COs coordinated and handled the administration of the WWD project. They began the initiative to build WWD and developed and iterated the data collection process, front-end WWD project website, back-end database, and data processing pipeline in accordance with community feedback. COs ensured accessibility was a top priority and continuously integrated feedback from the community to achieve this goal. Additionally, they helped to review community contributions for dishes from their regions. All COs also served as CCs and CAs for their home regions.
Together, the three stakeholder groups made collecting fine-grained regional dish data possible. The CAs played a crucial role in mediating interactions between CCs and the overall research project. Because CAs were members of the communities for which they served as ambassadors, they had intimate knowledge of the potential barriers to participation that CCs might face. For example, when the data collection instrument for WWD was being developed, CAs raised the need to build a web interface that was compatible with the 2G cellular networks that were the norm for Internet connectivity in their home regions. In the following section, we describe the technical system architecture through which data contributions were made and conclude with an overview of how community members interacted with and shaped the design of this system.
@@ 4.3.1. Data collection
Data collection occurred through a web-based form accessible on our website (see Figure B.1). On the web-based form, CCs were first provided with information about the project and a consent form. If they provided consent and indicated that they were over 18, they were then taken to the data collection form and asked to provide information about their chosen dish. As part of this form, CCs were asked to upload a non-generated image of food from personal records, indicate the dish name in its local language with a translation or phonetic equivalent to English when possible, and provide information about the customs associated with the dish. Please see Table 1 above for a full list of the fields CCs were asked to provide.
For the WWD data collection process to be useful, contributors had to be recruited and actively engaged. To that end, the COs and CAs worked together to ensure the data collection system was easily shareable via digital social networks to support multiple browsers, and also had interactive elements to amplify CC excitement and motivation. Recruitment involved the use of many social networks, including through existing communities such as the AYA Project, [^6] https://aya.for.ai/ Masakhane, [^7] https://www.masakhane.io/ the Deep Learning IndabaX network, [^8] https://deeplearningindaba.com/2024/indabax/ AI Saturdays Lagos, [^9] https://aisaturdayslagos.github.io/ and OLS. [^10] https://we-are-ols.org/ Posts were also placed on a mailing list to encourage engagement.
Engagement was explicitly encouraged through the promotion of a submission leaderboard (see Figure 3), which displayed the number of dish submissions and contributors received from each country or region. This was done in an attempt to gamify the experience and build excitement and fun. The leaderboard was promoted during recruitment outreach.
Figure 3. Screenshot of the World Wide Dishes Leaderboard showing the top 10 countries by number of dishes contributed, the total number of dishes contributed, and the total number of contributors per country.
@@ 4.3.2. Data storage
The collected data is stored in two sections: the dish data, which has been made public and is completely anonymised, and the personal data, which has been collected and stored according to the terms agreed by our ethics review board. We collected names, ages, and national identities, but also accepted anonymous submissions. Names were only used when explicit consent was given in order to acknowledge contributions publicly. Age and national identity data were required to help us understand the demographics of people submitting to WWD. Approval for data collection and the subsequent research study was obtained from the Departmental Research Ethics Committees of the Computer Science Department at Oxford University (reference: CS_C1A_24_004). Notably, the ethics review required us to collect age data, as CCs had to be over the age of 18 to participate; age data was therefore collected even if the CC otherwise submitted dish data anonymously.
@@ 4.3.3. Data processing
    Inconsistencies in submitted data were handled by consulting CAs where possible, and CAs also consulted with CCs as needed. However, we aimed to collect lived experiences around the world and so we did not make any efforts to police a ‘ground truth’ for each dish, or assign an ‘origin’ or ‘authenticity’ to any dish. Multiple nations or cultural groups can share the same dishes, for a wide variety of reasons including historical trade routes, war and occupation, the redrafting of political borders without regard to cultural borders, and migration patterns. Our task was not to make judgments about cultural ‘ownership’ of the submitted dishes. Instead, the data collection asked participants to offer dishes from their own lived experiences and personal backgrounds, so we expected to see similarities across borders, as well as some reasonable regional and preference variations of the ingredients of dishes within and across borders.
@@ 4.3.4. How stakeholders interacted with and shaped the design of the WWD technical system architecture
All three stakeholder groups (CCs, CAs, and COs) interacted with and shaped the technical system through which data contributions were made. Stakeholders surfaced additional metadata fields that were important to understanding the traditions of their local dishes during focus groups prior to data collection, and the COs then adjusted the data collection form to include these fields. By guiding the questions asked on the data collection form, community members shaped both the technical data collection system and the resulting WWD dataset. Moreover, COs and CAs who had intimate knowledge of potential barriers to participation for their home regions directly shaped the technical data collection system by advocating for a web interface that was mobile-first and compatible with low-bandwidth cellular networks. CAs provided valuable insight during data collection as to sub-regions and cultures that were underrepresented during data collection and COs and CAs responded by introducing extra efforts to reach these regions.
@@ 5. Methods and Analysis
Our study aims to map out the process of building a dataset of cultural objects (specifically, food) through a bottom-up, community-led approach. In this section, we provide a clear timeline for developing WWD, detail our data collection and reflective practices, and explain our analysis methodology.
@@ 5.1. Timeline and project context
Planning, building, and deploying the system that resulted in the WWD dataset was a process that occurred over 12 months. In October 2023, the first author initiated the WWD project. From November to December 2023, the team expanded to include three more COs, with four more joining in March 2024 (see Table 2 for details about the full team list). The final addition of COs occurred in April 2025 in order to include two CAs in honour of their exceptional commitment to the project. Between January and February 2024, the CO team developed the data collection instrument and website. The WWD website went live in February 2024. Inspired by the AYA initiative’s use of community ambassadors (Singh et al., 2024), the COs recruited and onboarded CAs throughout March 2024 to support community engagement. In April 2024, the COs and CAs conducted a month-long campaign for data contributions, resulting in 765 dishes submitted from 106 regions. The research activities conducted as part of the development of WWD were approved by the first author’s ethics review panel.